smithm29:

A 41-gram sample of a substance that's used to detect explosives has a k-value of 0.139. Find the half-life in days.

Vocaloid:

same formula as before N = N_0 ( e^(-kt)) N_0 = 41 N = 41/2 = 20.5 k = 0.139

Vocaloid:

N = N_0 ( e^(-kt)) 20.5 = 41 * e^(-0.139)t 0.5 = e^(-0.139)t ln(0.5) = -0.139t

smithm29:

so... 20.5=41^0.139t

Vocaloid:

yeah just don't forget the e

smithm29:

whats the e?

Vocaloid:

e is a constant that is (approximately) equal to 2.71828

smithm29:

oh ok

Vocaloid:

but you don't need to worry about it since it gets eliminated

Vocaloid:

taking the natural log will "undo" the e

smithm29:

cool

Vocaloid:

so if we have: ln(0.5) = -0.139t t = ln(0.5)/-0.139 = 5.0 days (depending on how they want you to round it)

Vocaloid:

my calc gives me 4.98667036374

smithm29:

days?

Vocaloid:

I think so

smithm29:

y does that look so easy... then u try it and its not

Vocaloid:

hm

smithm29:

no its just hard for me to learn something new with my ADD

Vocaloid:

ah, I see

smithm29:

wait im tryin to do one bms... after i divide both sides by N then what

Vocaloid:

you need to divide both sides by N_0 not N

Vocaloid:

you should get 0.5 on the left side and e(-something) on the right side

Vocaloid:

then you have: 0.5 = e(-kt) take the natural log of both sides

smithm29:

ok i got it

smithm29:

how do i get rid of the e?

Vocaloid:

when you take the natural log you can eliminate e since ln(e) = 1

Vocaloid:



smithm29:

so it would be N=-kt?

Vocaloid:

the ln(0.5) is still there

Vocaloid:

ln(0.5) = -kt

smithm29:

oh ok

Vocaloid:

then you just divide both sides by -k to get t

Vocaloid:

t = ln(0.5)/k, then just plug into your calculator

Vocaloid:

if the question asks for k instead it would be k = -ln(0.5)/t

smithm29:

what about the negative before the k?

Vocaloid:

you can just divide both sides by -1, which will "move" the negative sign over

Vocaloid:

ln(0.5) = -kt divide both sides by -1 gives -ln(0.5) = kt

Vocaloid:

** I made a typo earlier

smithm29:

so u still divide by -k?

Vocaloid:

yeah

Vocaloid:

ln(0.5) = -kt dividing by -k gives us -ln(0.5)/k = t

smithm29:

how do i get rid of the negative sign bc i dont think it would be -...days

Vocaloid:

ln(0.5) is also negative, so -ln(0.5) is positive

Vocaloid:

-ln(0.5) = negative * negative = positive

smithm29:

oh duh lol

smithm29:

im gonna close this and do my other question can u just tell me if im doin it right?

Vocaloid:

sure

smithm29:

thx